OpenStudy (anonymous):

Can someone check this for me. The square of a number increased by 5 times the number is equal to 176. Find the number. The number is 151 x+5^2=176 x=151

OpenStudy (welshfella):

x^2 + 5x = 176

OpenStudy (anonymous):

will x be the same on both sides

OpenStudy (welshfella):

not sure what you mean you need to solve the equation x^2 + 5x - 176 = 0 this is called a quadratic equation have you seen these before?

OpenStudy (anonymous):

yes i have i have trouble remembering #s can you give me a min

OpenStudy (anonymous):

x=11

OpenStudy (welshfella):

yes thats one answer whats the other?

OpenStudy (anonymous):

-16

OpenStudy (anonymous):

x=-32/2

OpenStudy (welshfella):

yes thats correct - i got in a tangle then

OpenStudy (anonymous):

a what?

OpenStudy (welshfella):

I had a brainstorm!

OpenStudy (anonymous):

The square of a number increased by 5 times the number is equal to 176 is -16

OpenStudy (anonymous):

I can only write one answer

OpenStudy (welshfella):

well write 11 then

OpenStudy (anonymous):

is this another brainstorm??? :)

OpenStudy (welshfella):

no lol

OpenStudy (anonymous):

ok thanks so my final answer is 11 Thank you

OpenStudy (welshfella):

11 fits perfectly:- 11^2 + 5*11 = 121 + 55 = 176

OpenStudy (anonymous):

I only have about 10 more review pages to go before my exam. omgosh :(

OpenStudy (welshfella):

good luck

OpenStudy (anonymous):

thanks:)

OpenStudy (welshfella):

yw and thank you - i get 75 for problem solving so i'm promoted to Mentor!!

OpenStudy (anonymous):

WooHoo so your prize is you get to correct my questions all night. :)

OpenStudy (welshfella):

lol!!

OpenStudy (welshfella):

sorry I have to go - Its pretty late here in Europe